The height if a parallelogram is 5 times its base. If the area of the parallelogram is 1805 square millimeters, find the length if its base and height. Any help would be much appreciated!!
Answer by stanbon(75887)

The height of a parallelogram is 5 times its base. If the area of the parallelogram is 1805 square millimeters, find the length if its base and height.
--------------------
Area = base*height
-----
1805 = b*5b
----
361 = b^2
----
base = 19 ml
----
height = 5x = 5*19 = 95 ml
-------------------
Cheers,
Stan H.
--------------------